2-x/3=x/2 answer is said to be 12 however i have to show my work and i have no idea how the answer became 12. i do have other problems but if this one can be explained then i will be able to do the rest on my own.
Answer by stanbon(75887)

2-x/3=x/2
Multiply thru by 6 to get:
12-2x = 3x
12 = 5x
x = 12/5
---------------
Checking:
2 - (12/5)/3 = (12/5)/2
2 - 12/15 = 12/10
(30/15) - (12/15) = 12/10
18/15 = 12/10
6/5 = 6/5
=============
Cheers,
Stan H.